# Unveiling Linguistic Regions in Large Language Models

### v2

@@ 5 Conclusion
This paper explores the pivotal role of certain parameters in Large Language Models (LLMs), identifying a core region essential for multilingual alignment and generalization. Removing this region causes a complete loss of linguistic ability in LLMs. Further more, we discover that this core region is concentrated in specific dimensions, perturbing only one dimension can cause a significant decrease in language ability. Moreover, beyond the core linguistic regions, we observe that monolingual regions exist within LLMs that affect specific languages. Importantly, we note that the catastrophic forgetting phenomenon during further pre-training may be related to drastic changes in core linguistic regions, as freezing this part during further pre-training alleviates the issue substantially. Our analysis and findings provide new perspectives and explanations for LLMs’ linguistic competence.
@@ Appendix B Core Linguistic Region
In Figure 9, we present the distribution of the ‘Top’ 5% regions of the LLaMA-2-13B model on the Attn.o matrix. The results indicate that across various layers, the core linguistic region on Attn.o matrix is concentrated on different rows. This difference is observed among the 40 attention heads in the 13B model.
@@ Appendix D Single Parameter Perturbation
In a Transformer block, each column in the Attn.o and the MLP.down matrix of the feed-forward layer can be considered as the input weights of a neuron. Thus, perturbing a column can be seen as disturbing the input weights of a neuron. Viewed from another angle, if we disturb the output activation value of this neuron, a similar effect should be observed. Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight. Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons. In Figure 7 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix.
@@ Appendix E Ablation Study
| Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) | 1 | 848.326 | 27.265 |
@@ Appendix F Monolingual Region
In Figure 8, we use the prompt "There are 365 days in a year and 12" to test the model’s output in English, Arabic, and Chinese, respectively. The results indicate that removing the monolingual regions causes the model to lose the relative language competence, leading it to generate repetitive, nonsensical responses rather than correct answers like "12 months in a year".

### v3

@@ 5 Conclusion
This paper explores the pivotal role of certain parameters in Large Language Models (LLMs), identifying a core region essential for multilingual alignment and generalization. Removing this region causes a complete loss of linguistic competence in LLMs. Furthermore, we discover that this core region is concentrated in specific dimensions, perturbing only one dimension can cause a significant decrease in linguistic ability. Moreover, beyond the core linguistic regions, we observe that monolingual regions exist within LLMs that affect specific languages. Importantly, we note that the catastrophic forgetting phenomenon during further pre-training may be related to drastic changes in core linguistic regions, as freezing this part during further pre-training alleviates the issue substantially. Our analysis and findings provide new perspectives and explanations for LLMs’ linguistic competence.
@@ Appendix B Core Linguistic Region
In Figure 9, we present the distribution of the ‘Top’ 5% regions in the Attn.o matrix for the LLaMA-2-13B model. The results indicate that across various layers, the core linguistic region on Attn.o matrix is concentrated on different rows. This difference is observed among the 40 various attention heads.
@@ Appendix D Single Parameter Perturbation
In a Transformer block, each column in the Attn.o and the MLP.down matrix of the FFN layer can be considered as the input weights of a neuron. Thus, perturbing a column can be seen as disturbing the input weights of a neuron. Viewed from another angle, if we disturb the output activation value of this neuron, a similar effect should be observed. Within LLaMA, there is a specific module called RMSNorm, where each dimension is associated with a weight. Perturbations to these weights can be regarded as disturbances to the output activation values of the corresponding neurons. In Figure 7 (right), we visually demonstrate how RMSNorm affects a column of the Attn.o and the FFN.down matrix.
@@ Appendix E Ablation Study
| Attn.o(row) Attn.k/q/v(column) FFN.down(column) | 1 | 848.326 | 27.265 |
@@ Appendix F Monolingual Region
In Figure 8, we use the prompt "There are 365 days in a year and 12" to test the model’s output in English, Arabic, and Chinese, respectively. The results indicate that removing the monolingual regions causes the model to lose the relative language competence, leading the model to generate repetitive, nonsensical responses rather than correct answers like "12 months in a year".